# Non-Invasive Medical Digital Twins using Physics-Informed Self-Supervised Learning

### v1

@@ 6 Conclusion
We present methodology for non-invasive prediction of patient specific physics-based models directly from imaging that can form the basis for further development of medical digital twins. Our experiments showcase the ability of these methods to personalize a high fidelity physics-based model from video data. We demonstrate that information beyond labels is learned in the setting of echocardiography for cardiac pressure-volume loops. Our work may be extended to improve mathematical modeling for medical or non-medical applications. Latent parameters in our experimental setting are clinically informative, and our methodology gives a non-invasive process for estimating their value. Measurements of cardiac valve resistance in the clinical setting are typically obtained by taking auxiliary measurements from invasive catheters and plugging into mathematical formulas (Mascherbauer et al., 2004). We predict resistance as latent parameters from echocardiography. Further development of our model and methodology could contribute new non-invasive direct ways of computing such parameters.
With further development, our model can provide the basis for in-silico digital twins studies that do not require expensive or invasive set-up. Theoretically, given complete knowledge of the patient states $\mathbf{x}$, the full set of parameters of our mathematical are identifiable (Appendix B). In practice, our medical image inputs display only partial knowledge of patient states and computational techniques are insufficient. Reducing the number of learned parameters without comprising model generality might improve the ability of our methods to distinguish non-labeled states. Future work could explore methods to reduce model complexity before training generative solutions (see Appendix D.2).
Impact Statement
This paper presents work contributing a novel theoretical approach that can be further iterated and tested toward the development of personalized medical tools for eventual use in clinical settings. The increasing development of tools and theoretical methodology for precision medicine aims has the potential to have widespread downstream positive consequences for access and accuracy in patient care. However, careful vetting and testing are critical in this setting, as machine learning for medicine can very easily cause great harm with small inaccuracies.
Our experiments utilized de-identified data, previously made publicly available, adhering to ethical considerations.
Acknowledgements
The authors would like to thank Ellen Roche, Caglar Ozturk, Yiling Fan and Luca Rosalia for the helpful discussions.
This material is based upon work supported by the National Science Foundation Graduate Research Fellowship under Grant No. DGE 2146752.
The code is provided at https://github.com/AlaaLab/CardioPINN for reproducibility and further exploration.
Appendix A Mathematical formulation of lumped-parameter hydraulic analogy model (Simaan et al., 2008)
The (Simaan et al., 2008) model of the cardiovascular system is characterized by the values of the circuit elements, including the capacitances, resistances, and inductance, and by the initial conditions. The model is governed by 17 parameters: eight static parameters that account for the circuit’s fixed properties, four parameters that capture the elastance function, and five values describing the initial conditions (see Appendix A.2). The elastance function relates the LV volume to the LV pressure
@@ Acknowledgements
where $V_{d}$ is the patient specific theoretical volume of the LV at zero pressure. The elastance function is modeled in (Simaan et al., 2008) as
@@
where $P_{LV},P_{LA},P_{A},P_{Ao}$ are pressures of the left-ventricle, left atrium, arteries, and aorta, respectively, and $Q_{T}$ is total blood flow. We perform a change of variables from the (Simaan et al., 2008) formulation of the model and write
@@
  • For $P_{LA}(0)$, the initial pressure in the left atrium, we know that the filling phase has finished, and the pressures in the left atrium and LV must be equal (Sainte-Marie et al., 2006).
@@ A.1 Addition of a feedback-controlled LVAD
We simulate the addition of a left ventricular assist device (LVAD) to the cardiac system. One end of the device is connected between the left ventricle and the aorta, and the other end is connected between the left ventricle and the left atrium. Many different LVADs have been developed; we take here the case of a rotary-controlled LVAD as described in (Choi et al., 1997; Simaan et al., 2008).
The presence of the LVAD induces the addition of a new state variable, $x_{6}$, representing the blood flow through the device. In addition, the LVAD representation in the electric circuit is described by parameters $R_{0}$, $R_{i}$, $\bar{P}$, $\alpha$, $L_{i}$, $L_{0}$, $\beta_{0}$, $\beta_{1}$, $\beta_{2}$, — describing compliances, static resistances and a time-varying resistances— and by the pump rotational speed, $\omega(t)$, which is a function of time. The pump speed can be defined beforehand (e.g. a linear increasing function), or it can be updated in real-time based on the state of the system. A more detailed description of the parameters and equations of the LVAD can be found elsewhere (Choi et al., 1997; Simaan et al., 2008).
@@
for new state vector $\mathbf{y}$, containing the first five variables as in 19 in addition to the blood flow through the LVAD, $x_{6}(t)$. The new matrices of the system are
@@
The vector $p(\mathbf{x})$ is as defined in 23, and $B(t)$ is given by
@@
A.2 Parameters of the lumped-parameter hydraulic analogy model (Simaan et al., 2008)
@@
$R_{C}$
Characteristic resistance
[…]
$R_{S}$
Systemic vascular resistance, related to the level of activity of the patient (Ferreira et al., 2005)
[…]
$C_{A}$
Aortic compliance
0.0800 ml/mmHg (Simaan et al., 2008)
Fixed (0.08 ml/mmHg)
$C_{S}$
Systemic compliance
1.3300 ml/mmHg (Simaan et al., 2008; Yu, 1998)
Fixed (1.33 ml/mmHg)
$C_{R}$
Left atrial compliance, represents preload and pulmonary circulation (Simaan et al., 2008)
4.4000 ml/mmHg (Simaan et al., 2008; Yu, 1998)
Fixed (4.4 ml/mmHg)
$L_{S}$
Inertance of blood in the aorta
[…]
@@
$E_{MAX}$
Maximum elastance, related to the contractility of the heart (Stergiopulos et al., 1996)
1.5-2.0 mmHg/ml (Simaan et al., 2008), 2.31 (Stergiopulos et al., 1996), 0.7-2.5 mmHg/ml (Ferreira et al., 2005)
[0.5, 3.5] mmHg/ml
$E_{MIN}$
Minimum elastance
0.06 mmHg/ml (Stergiopulos et al., 1996), 0.05 mmHg/ml (Simaan et al., 2008)
[0.02, 0.1] mmHg/ml
$V_{d}$
Theoretical LV volume at zero pressure
10 ml (Ferreira et al., 2005), 12 ml (Simaan et al., 2008), 20 ml (Stergiopulos et al., 1996)
[4.0, 25.0] ml
$T_{c}$
Heart cycle duration ($T_{c}=60/HR$)
0.8-1.0 s (Simaan et al., 2008)
[0.4, 1.7] s (i.e. 35-150 beats/min)
@@
start_v
Initial condition related (and not necessarily equal) to the ED LV volume
140 ml (Simaan et al., 2008)
[0, 280.0] ml
start_pao
Initial condition related (and not necessarily equal) to the ED aortic pressure
$\sim 77$ mmHg (Simaan et al., 2008)
Fixed (75 mmHg)
@@
$R_{o}$
Outlet resistance of cannulae
0.0677 mmHg.s/ml (Simaan et al., 2008)
Fixed (0.0677 mmHg.s/ml)
$R_{i}$
Inlet resistance of cannulae
0.0677 mmHg.s/ml (Simaan et al., 2008)
Fixed (0.0677 mmHg.s/ml)
$\alpha$
LVAD pressure parameter
-3.5s/ml (Simaan et al., 2008)
Fixed (-3.5s/ml)
$\bar{P}$
LVAD weight parameter
1 mmHg (Simaan et al., 2008)
Fixed (1 mmHg)
$L_{i}$
Inlet inertance of cannulae
0.0127 mmHg.s/ml (Simaan et al., 2008)
Fixed (0.0127 mmHg.s/ml)
$L_{o}$
Outlet inertance of cannulae
0.0127 mmHg.s/ml (Simaan et al., 2008)
Fixed (0.0127 mmHg.s/ml)
$\beta_{0}$
LVAD dependent pressure parameter
-0.296 (Choi et al., 1997)
Fixed (-0.296)
$\beta_{1}$
LVAD dependent pressure parameter
-0.027 (Choi et al., 1997)
Fixed (-0.027)
$\beta_{2}$
LVAD dependent pressure parameter
$9.9025\times 10^{-7}$ (Simaan et al., 2008)
Fixed ($9.9025\times 10^{-7}$)
$H$
Circuit pressure difference (inlet-outlet) defining LVAD pressure parameters
$H=\beta_{0}x_{6}+\beta_{1}\frac{dx_{6}}{dt}+\beta_{2}\omega^{2}$
Appendix B Identifiability of the ODE inverse problem $\mathcal{M}$
We consider the inverse problem given by the (Simaan et al., 2008) model $\mathcal{M}(\theta)=\mathbf{x}$. We argue that this function is well defined: given $\theta$, which includes parameters describing the initial system state, there is a unique ODE solution $\mathbf{x}$. Using Picard-Lindelöf, it is sufficient for $A_{1}\mathbf{x}+D_{1}p(\mathbf{x})$ to be continuous in $t$ and Lipschitz continuous in $\mathbf{x}$. As $\mathbf{x}$ describes pressures and flows that cannot be discontinous in $t$, the first condition is satisfied. As $A_{1}\mathbf{x}+D_{1}p(\mathbf{x})$ is linear in each of the four cases defined by the pairwise possibilities for the rows of $p(\mathbf{x})$ to be either positive or zero, we can take the maximum needed Lipschitz constant amongst these four possibilities to see the entire equation is Lipschitz continuous.
@@ Appendix B Identifiability of the ODE inverse problem $\mathcal{M}$
The vector $L(p(\mathbf{x}))$ is constant. For any fixed $s$, $L(x)(s)$ is also a constant. Thus, each $s$ gives a linear system of five equations in the eight unknown static parameters governing the (Simaan et al., 2008) model. Identifying these eight parameters uniquely simply requires choosing enough values of $s$ to solve each row as its own linear system. This shows that $\mathcal{M}$ is injective.
@@ C.1 Ill-posed inverse problems
Generally, inverse problems are the task of determine inputs $\theta\in\Theta$ from measurements $\mathbf{y}\in\mathbf{Y}$ in the set up $\mathcal{F}(\theta)=\mathbf{y}$ with forward operator $\mathcal{F}:\Theta\to\mathbf{Y}.$
Inverse problems are typically ill-posed, meaning that the forward model $\mathcal{F}$ may be non-injective or that small perturbations of measurements $\mathbf{y}$ (i.e. noise) cause large change in the corresponding inputs $\theta$. Often, a variational approach is taken to mitigate these challenges, meaning the problem is reformulated into the task of minimizing over the possible solution set $\Theta$ the sum of a data fidelity term $D$ to represent fitting the model $\mathcal{F}$ in addition to a term incorporating solution criteria $R$ at some level $\alpha$:
@@
Paired $\{(\mathbf{y}_{i},\theta_{i})\}$
Unpaired $\{\mathbf{y}_{j}\},\{\theta_{i}\}$
Only $\{\theta_{i}\}$
Only $\{\mathbf{y}_{i}\}$
Paired $\{(\mathbf{y}_{i},\mathbf{x}_{i})\}$
No or limited data
Known $\mathcal{F}$
Reconstruction techniques (Jin et al., 2017), Denoising autoencoder (Mousavi et al., 2015)
Same as training with paired $(\theta,\mathbf{y})$
Same as training with paired $(\theta,\mathbf{y})$
UPINN (Dubois et al., 2022)
Traditional PINNs for Equation Discovery (Raissi et al., 2019)
Unknown $\mathcal{F}$
Supervised learning (Zhu et al., 2018)
With data distributions: (del Álamo, 2023), CycleGAN (Zhu et al., 2017)
Know distribution for $\mathcal{F}$
Blind demodulation (Hand & Joshi, 2019), DeblurGAN (Kupyn et al., 2018)
AmbientGAN (Bora et al., 2018)
Know $\mathcal{M}$
Our method
@@ C.2 Physics informed neural networks and related approaches
In big data settings, physical laws may be learned during modeling. In smaller data settings, incorporating physics into models can improve performance (von Rueden et al., 2023). This is especially pertinent to personalized medicine where data is inherently limited. PINNs typically have the goal of either finding solutions (forward) or system parameters (inverse) to systems of differential equations (Raissi et al., 2019; Cuomo et al., 2022). PINN methodologies define training loss as the sum of a data loss and a regularization term, which is the differential equation itself. There is extensive development and mixed success in PINN models for both goals (Krishnapriyan et al., 2021).
@@
Efforts to incorporate knowledge into neural network models to improve training and accuracy include techniques to leverage physical models represented by differential equations (von Rueden et al., 2023). In big data settings, physical laws may be learned during modeling. In smaller data settings, incorporating physics into models can improve performance (von Rueden et al., 2023). This is especially pertinent to personalized medicine where data is inherently limited. Physics-informed neural networks (PINNs) typically have the goal of either finding solutions (forward) or system parameters (inverse) to systems of differential equations (Raissi et al., 2019; Cuomo et al., 2022). PINN methodologies define training loss as the sum of a data loss and a regularization term, which is the differential equation itself. There is extensive development and mixed success in PINN models for both goals (Krishnapriyan et al., 2021).
@@
The set up is the following. Given a state vector $\mathbf{x}\in\mathbf{X}$ dependent on time variable $t$ and differential equation $f(\theta,\mathbf{x})=0$ constraining the solutions $\mathbf{x}$ subject to a parametrization $\theta$ with initial or boundary data $D(\theta,\mathbf{x})$, PINNs construct a variational minimization task
$\min D(\theta,\mathbf{x})+f(\theta,\mathbf{x})$ (30)
@@
so that $f$ acts as a regularizer. This framework can be used to solve both the forward task of solving $f$ for $\mathbf{x}$ and the inverse task of finding $\theta$. The minimization is over the parameters of the neural network in the case of the forward task and over $\theta\in\Theta$ in the inverse task. We call the forward operator in this set up $\mathcal{M}$, which is the task of solving the differential equation.
@@
Drawbacks of the PINN approach include that the differential equation and boundary conditions need not be exactly solved. Other approaches have solved $f$ as a hard constraint for these same aims in physics-constrained learning (PCL), which mitigates this challenge. A variety of variants on these traditional approaches continue to develop for both forward and inverse problems in various specific problem set ups such as fPINNs (Pang et al., 2019), SPINNs (Ramabathiran & Ramachandran, 2021), B-PINNs (Yang et al., 2021), VPINNs (Kharazmi et al., 2021), and cPINNs (Jagtap et al., 2020) which while not directly relevant in our experiments may be valuable in future digital twin tasks. We compare related techniques to our own with examples in Table 4.
@@
Forward task (learn $\mathbf{x}(t)$)
Inverse task (learn $\theta$)
Traditional PINNs with $f$ as regularization term
(Kissas et al., 2020) (Herrero Martin et al., 2022) (Sahli Costabal et al., 2020) (Grandits et al., 2021)
(van Herten et al., 2022)(Kissas et al., 2020)(Herrero Martin et al., 2022)
PCL: $f$ as hard constraint
(Zhao et al., 2019)
(Xu & Darve, 2022)
PI-NODEs: Train a neural network to learn the non-linear dynamics and a Neural ODE (Chen et al., 2018) to learn linear components
(Lai et al., 2021)
(Lai et al., 2021)
P-SSL: $f$ is used to generate synthetic data and use data loss for training a generative solution to inform a larger inverse task
Our method
Our method could be inverted for this task
Neural ODEs
(Chen et al., 2018) proposed the use of differential equation solvers for solving neural networks upon noticing the similarity between ResNet gradient descent and Euler’s method. Viewing neural networks as ODEs means forward propagation is the system solution. Our fixed pretraining task for leveraging the physical model shares similarities with Neural ODEs in that we train a network to learn to solve an ODE. Our aim differs as forward propagation instead maps latent parameters to state variables.
Neural ODEs are trained by back propagating through a numerical ODE solver, which would have been a direct way to generate $\mathcal{M}$ alternative to our pretraining approach. The computation cost of this strategy was prohibitive in our case. Other strategies for reducing computational costs associated to back propagating through ODE solvers show up in Neural ODE methodology include adjoint state variable methods (Chen et al., 2018) and augmenting the system to a higher dimensional space (Dupont et al., 2019). Our synthetic data-driven pretraining approach further reduces computational costs and improves model convergence.
Appendix D Electric circuit derivation of the (Simaan et al., 2008) model

### v2

@@ 5 Conclusion
We present methodology, Med-Real2Sim, for non-invasive prediction of patient specific physics-based models directly from imaging that can form the basis for further development of medical digital twins. Our experiments showcase the ability of these methods to personalize a high fidelity physics-based model from video data. We demonstrate that information beyond labels is learned in the setting of echocardiography for cardiac pressure-volume loops. Our work may be extended to improve mathematical modeling for medical or non-medical applications. Latent parameters in our experimental setting are clinically informative, and our methodology gives a non-invasive process for estimating their value. Further development of our methods could contribute new non-invasive direct ways of computing such parameters and provide the basis for in-silico digital twin studies.
Limitations We highlight that our model experiments are limited in their clinical validation. In using only publicly available datasets with limited labels, we are not able to compare measurements of heart resistance or left-ventricular pressure, which would be necessary to fine-tune such a model for clinical use. While the ability to predict latent states and parameters is the core novelty of our research and approach, acquiring further data to on these latent states could improve the model substantially. This is a critical future direction for work in building digital twins with our methodology.
Appendix A Mathematical formulation of lumped-parameter hydraulic analogy model
The Simaan et al. (65) model of the cardiovascular system is characterized by the values of the circuit elements, including the capacitances, resistances, and inductance, and by the initial conditions. The model is governed by 17 parameters: eight static parameters that account for the circuit’s fixed properties, four parameters that capture the elastance function, and five values describing the initial conditions (see Appendix A.2). The elastance function relates the LV volume to the LV pressure
@@ Appendix A Mathematical formulation of lumped-parameter hydraulic analogy model
where $V_{d}$ is the patient specific theoretical volume of the LV at zero pressure. The elastance function is modeled in (65) as
@@
where $P_{LV},P_{LA},P_{A},P_{Ao}$ are pressures of the left-ventricle, left atrium, arteries, and aorta, respectively, and $Q_{T}$ is total blood flow. We perform a change of variables from the (65) formulation of the model and write
@@
  • For $P_{LA}(0)$, the initial pressure in the left atrium, we know that the filling phase has finished, and the pressures in the left atrium and LV must be equal (62).
@@ A.1 Addition of a feedback-controlled LVAD
We simulate the addition of a left ventricular assist device (LVAD) to the cardiac system. One end of the device is connected between the left ventricle and the aorta, and the other end is connected between the left ventricle and the left atrium. Many different LVADs have been developed; we take here the case of a rotary-controlled LVAD as described in (12; 65).
The presence of the LVAD induces the addition of a new state variable, $x_{6}$, representing the blood flow through the device. In addition, the LVAD representation in the electric circuit is described by parameters $R_{0}$, $R_{i}$, $\bar{P}$, $\alpha$, $L_{i}$, $L_{0}$, $\beta_{0}$, $\beta_{1}$, $\beta_{2}$, — describing compliances, static resistances and a time-varying resistances— and by the pump rotational speed, $\omega(t)$, which is a function of time. The pump speed can be defined beforehand (e.g. a linear increasing function), or it can be updated in real-time based on the state of the system. A more detailed description of the parameters and equations of the LVAD can be found elsewhere (12; 65).
@@
for new state vector $\mathbf{y}$, containing the first five variables as in 18 in addition to the blood flow through the LVAD, $x_{6}(t)$. The new matrices of the system are
@@
The vector $p(\mathbf{x})$ is as defined in 22, and $B(t)$ is given by
@@
A.2 Parameters of the lumped-parameter hydraulic analogy model
@@
$C_{A}$ Aortic compliance 0.0800 ml/mmHg (65) Fixed (0.08 ml/mmHg)
$C_{S}$ Systemic compliance 1.3300 ml/mmHg (65; 82) Fixed (1.33 ml/mmHg)
$C_{R}$ Left atrial compliance, represents preload and pulmonary circulation (65) 4.4000 ml/mmHg (65; 82) Fixed (4.4 ml/mmHg)
@@
$E_{MAX}$ Maximum elastance, related to the contractility of the heart (66) 1.5-2.0 mmHg/ml (65), 2.31 (66), 0.7-2.5 mmHg/ml (22) [0.5, 3.5] mmHg/ml
$E_{MIN}$ Minimum elastance 0.06 mmHg/ml (66), 0.05 mmHg/ml (65) [0.02, 0.1] mmHg/ml
$V_{d}$ Theoretical LV volume at zero pressure 10 ml (22), 12 ml (65), 20 ml (66) [4.0, 25.0] ml
$T_{c}$ Heart cycle duration ($T_{c}=60/HR$) 0.8-1.0 s (65) [0.4, 1.7] s (i.e. 35-150 beats/min)
@@
start_v Initial condition related (and not necessarily equal) to the ED LV volume 140 ml (65) [0, 280.0] ml
start_pao Initial condition related (and not necessarily equal) to the ED aortic pressure $\sim 77$ mmHg (65) Fixed (75 mmHg)
@@
$R_{o}$ Outlet resistance of cannulae 0.0677 mmHg.s/ml (65) Fixed (0.0677 mmHg.s/ml)
$R_{i}$ Inlet resistance of cannulae 0.0677 mmHg.s/ml (65) Fixed (0.0677 mmHg.s/ml)
$\alpha$ LVAD pressure parameter -3.5s/ml (65) Fixed (-3.5s/ml)
$\bar{P}$ LVAD weight parameter 1 mmHg (65) Fixed (1 mmHg)
$L_{i}$ Inlet inertance of cannulae 0.0127 mmHg.s/ml (65) Fixed (0.0127 mmHg.s/ml)
$L_{o}$ Outlet inertance of cannulae 0.0127 mmHg.s/ml (65) Fixed (0.0127 mmHg.s/ml)
$\beta_{0}$ LVAD dependent pressure parameter -0.296 (12) Fixed (-0.296)
$\beta_{1}$ LVAD dependent pressure parameter -0.027 (12) Fixed (-0.027)
$\beta_{2}$ LVAD dependent pressure parameter $9.9025\times 10^{-7}$ (65) Fixed ($9.9025\times 10^{-7}$)
$H$ Circuit pressure difference (inlet-outlet) defining LVAD pressure parameters $H=\beta_{0}x_{6}+\beta_{1}\frac{dx_{6}}{dt}+\beta_{2}\omega^{2}$
Appendix B Identifiability of the ODE inverse problem in the cardiac experiments
A unique and valuable attribute of the ordinary differential equation (ODE) in our cardiac hemodynamics experiments is its identifiability. In this experiment setting, we can demonstrate that the model $\mathcal{M}$ is well-defined and injective. To show this, we consider the inverse problem posed by finding parameters of the Simaan et al. (65) model, written $\mathcal{M}(\theta)=\mathbf{x}$.
First, this function is well defined: given $\theta$, which includes parameters describing the initial system state, there is a unique ODE solution $\mathbf{x}$. Using Picard-Lindelöf, it is sufficient for $A_{1}\mathbf{x}+D_{1}p(\mathbf{x})$ to be continuous in $t$ and Lipschitz continuous in $\mathbf{x}$. As $\mathbf{x}$ describes pressures and flows that cannot be discontinuous in $t$, the first condition is satisfied. As $A_{1}\mathbf{x}+D_{1}p(\mathbf{x})$ is linear in each of the four cases defined by the pairwise possibilities for the rows of $p(\mathbf{x})$ to be either positive or zero, we can take the maximum needed Lipschitz constant amongst these four possibilities to see the entire equation is Lipschitz continuous.
@@ Appendix B Identifiability of the ODE inverse problem in the cardiac experiments
The vector $L(p(\mathbf{x}))$ is constant. For any fixed $s$, $L(x)(s)$ is also a constant. Thus, each $s$ gives a linear system of five equations in the eight unknown static parameters governing the Simaan et al. (65) model. Identifying these eight parameters uniquely simply requires choosing enough values of $s$ to solve each row as its own linear system. This shows that $\mathcal{M}$ is injective.
@@ C.1 Ill-posed inverse problems
Generally, inverse problems are the task of determine inputs $\theta\in\Theta$ from measurements $\mathbf{y}\in\mathbf{Y}$ in the set up $\mathcal{F}(\theta)=\mathbf{y}$ with forward operator $\mathcal{F}:\Theta\to\mathbf{Y}.$ Inverse problems are typically ill-posed, meaning that the forward model $\mathcal{F}$ may be non-injective or that small perturbations of measurements $\mathbf{y}$ (i.e. noise) cause large change in the corresponding inputs $\theta$. Often, a variational approach is taken to mitigate these challenges, meaning the problem is reformulated into the task of minimizing over the possible solution set $\Theta$ the sum of a data fidelity term $D$ to represent fitting the model $\mathcal{F}$ in addition to a term incorporating solution criteria $R$ at some level $\alpha$:
@@
For example, $D$ might be $||\mathcal{F}(\theta)-\mathbf{y}||$ to encourage model fitting and $R$ might be $||\theta||$ to encourage input sparsity. This optimization task can been solved using analytical techniques to find $\theta$ directly or using generative learning techniques to approximate $\mathcal{F}^{-1}$ to find $\theta$.
@@
These existent set ups differ from our task as our partially known model is really the composition of a known and unknown model. We take a generative inverse problem approach and develop a model for $\mathcal{F}^{-1}$, which we compare to other generative solutions in Table 4 across the following two taxonomies: known versus unknown forward model $\mathcal{A}$ (row) and paired versus unpaired input-output data (column). This taxonomy mirrors that in (50).
Paired $\{(\mathbf{y}_{i},\theta_{i})\}$ Unpaired $\{\mathbf{y}_{j}\},\{\theta_{i}\}$ Only $\{\theta_{i}\}$ Only $\{\mathbf{y}_{i}\}$ Paired $\{(\mathbf{y}_{i},\mathbf{x}_{i})\}$ No or limited data
Known $\mathcal{F}$ Reconstruction techniques (33), Denoising autoencoder (48) Same as training with paired $(\theta,\mathbf{y})$ Same as training with paired $(\theta,\mathbf{y})$ UPINN (18) - Traditional PINNs for Equation Discovery (59)
Unknown $\mathcal{F}$ Supervised learning (84) With data distributions: (16), CycleGAN (85) - - - -
Know distribution for $\mathcal{F}$ Blind demodulation (27), DeblurGAN (40) AmbientGAN (5) - -
Know $\mathcal{M}$ Our method -
@@ C.2 Physics informed neural networks and related approaches
Efforts to incorporate knowledge into neural network models to improve training and accuracy include techniques to leverage physical models represented by differential equations (74). In big data settings, physical laws may be learned during modeling. In smaller data settings, incorporating physics into models can improve performance (74). This is especially pertinent to personalized medicine where data is inherently limited. Physics-informed neural networks (PINNs) typically have the goal of either finding solutions (forward) or system parameters (inverse) to systems of differential equations (59; 15). PINN methodologies define training loss as the sum of a data loss and a regularization term, which is the differential equation itself. There is extensive development and mixed success in PINN models for both goals (39).
@@
The set up is the following. Given a state vector $\mathbf{x}\in\mathbf{X}$ dependent on time variable $t$ and differential equation $f_{\theta}(\mathbf{x})=0$ constraining the solutions $\mathbf{x}$ subject to a parametrization $\theta$ with initial or boundary data $D(\theta,\mathbf{x})$, PINNs construct a variational minimization task
$\min D(\theta,\mathbf{x})+f_{\theta}(\mathbf{x})$ (29)
@@
so that $f_{\theta}(\mathbf{x})$ acts as a regularizer. This framework can be used to solve both the forward task of solving $f_{\theta}$ for $\mathbf{x}$ and the inverse task of finding $\theta$. The minimization is over the parameters of the neural network in the case of the forward task and over $\theta\in\Theta$ in the inverse task. We call the forward operator in this set up $\mathcal{M}$, which is the task of solving the differential equation.
Several papers have used PINN approaches to study cardiac function, especially blood flow, including papers that use non-invasive medical imaging data. van Herten et al. (70) learn parameters and solutions of a model for myocardial perfusion using measurements from MRIs. Kissas et al. (38) use a PINN approach to calibrate flow simulation model boundary conditions using data extracted from MRIs. Herrero Martin et al. (28) study cardiac electrophysiology with PINNs and demonstrate the ability of the method to learn a latent variable. Sahli Costabal et al. (61) and Grandits et al. (25) use a PINN approach to study cardiac activation. Our set-up differs in that we learn from images rather than through measurements obtained from images, and we predict informative latent parameters directly and seek a model that works for an entire patient population as opposed to a single patient. Our methodology differs in that the physical model is not integrated into the loss function during training. Instead, we use supervised learning trained on data loss alone to generate a forward model solving the ODE system and transfer this learning.
@@
Drawbacks of the PINN approach include that the differential equation and boundary conditions need not be exactly solved. Other approaches have solved $f_{\theta}$ as a hard constraint for these same aims in physics-constrained learning (PCL), which mitigates this challenge. A variety of variants on these traditional approaches continue to develop for both forward and inverse problems in various specific problem set ups such as fPINNs (53), SPINNs (60), B-PINNs (81), VPINNs (36), and cPINNs (32), which while not directly relevant in our experiments may be valuable in future digital twin tasks. We compare related techniques to our own with examples in Table 5.
Forward task (learn $\mathbf{x}(t)$) Inverse task (learn $\theta$)
Traditional PINNs with $f$ as regularization term (38) (28) (61) (25) (70)(38)(28)
PCL: $f$ as hard constraint (83) (80)
PI-NODEs: Train a neural network to learn the non-linear dynamics and a Neural ODE (10) to learn linear components (41) (41)
P-SSL: $f$ is used to generate synthetic data and use data loss for training a generative solution to inform a larger inverse task Our method Our method could be inverted for this task
@@
C.3 Neural ODEs
Chen et al. (10) proposed the use of differential equation solvers for solving neural networks upon noticing the similarity between ResNet gradient descent and Euler’s method. Viewing neural networks as ODEs means forward propagation is the system solution. It also means that, in contrast, to ResNet architectures, the network does not have a fixed number of layers. Neural ODEs are typically used to model time series data to replace ResNets. The set up is as follows.
Let ODESolver represent a numerical ordinary differential equation (ODE) solver. A fully connected neural network $f$ will represent the ODE the network learns to solve so that a forward pass looks like
$\mbox{ODESolver}(f,t_{0},x_{0},T)$ (30)
for initial condition $\mathbf{x}(t_{0})=x_{0}$ and prediction time series $T$. In this set up, $f$ can represent an ODE that we seek to approximate, analogous to our setting. Neural ODEs are trained by back propagating through the numerical ODE solver. Methods to reduce the computational cost of this backpropagation include the adjoint state variable methods in (10) and augmenting the system to a higher dimensional space (20).
Our fixed pre-training task for leveraging the physical model shares similarities with Neural ODEs in that we train a network to learn to solve an ODE. Our aim differs as a Neural ODE can learn the dynamics for a single patient. The ODE $f$ will be specific to a single patient parameter set $\theta$, but we seek to learn a map for a patient population.
Appendix D Electric circuit derivation of the (65) model